OpenStudy (anonymous):

A biologist is researching a newly-discovered species of bacteria. At time t = 0 hours, he puts two hundred bacteria into what he has determined to be a favorable growth medium. Eight hours later, he measures 1200 bacteria. Assuming exponential growth, what is the growth constant "k" for the bacteria? (Round k to three decimal places.

OpenStudy (zpupster):

Lets use this formula for exponential growth A= Pe^(kt) A= Amount of Population ending P = Population at start e= e k= growth constant t = time plugging in 1200 = 200e^(8k) divide both sides by 200 1200/200 =200e^(8k)/ 200 then take natural log both sides ln(6) = 8k divide 8 both sides use cal and you should get your answer

OpenStudy (anonymous):

@Princer_Jones i have no clue on how to solve this

OpenStudy (anonymous):

ok. so do you know what is meant by exponential growth??

OpenStudy (zpupster):

i showed all but the last steps ok divide both sides by 8 ln(6)/8 = 8k/8 k = Ln(6)/8 on calculator punch in 6 and there should be a button for natural log Ln ln6 = 1.79 1.79/8 = .22375/hr

OpenStudy (anonymous):

@_ashleeey

OpenStudy (zpupster):

ex growth is 2,4,8,16,32,64,128,256,512,1024,2048,4096.... or 2^2,2^3,2^4,2^5,2^6,2^7 the exponent grows and the growth doubles

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Exponential_growth

OpenStudy (anonymous):

@Princer_Jones no. i havent gotten the hang of this topic whatsoever

OpenStudy (anonymous):

what about this one A recent earthquake measured 6.8 on the Richter scale. How many times more intense was this earthquake than an earthquake that measured 4.7 on the Richter scale?

OpenStudy (anonymous):

did you red the link I gave? It will take only 5 mins and then you see the solution @zpupster gave... If you have problem, then tell me,,. @_ashleeey

OpenStudy (anonymous):

*read

OpenStudy (anonymous):

didnt see it, one sec.

OpenStudy (anonymous):

Just read the definiton, and then tell me.

OpenStudy (anonymous):

i read it

OpenStudy (anonymous):

ok so the formula is x(t)=x(0)*exp(kt) where k is the growth contant we need to calculate x(0) the amount in time 0 and x(t) the amount in time t

OpenStudy (anonymous):

@_ashleeey

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

all the data are already provided see x(0)=200 x(8)=1200 in the question

OpenStudy (anonymous):

@_ashleeey

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

so you can continue now i think, you just need to insert the values in the formula and calculate k

OpenStudy (anonymous):

@_ashleeey

OpenStudy (anonymous):

great, ill get to it, do you still have time?

OpenStudy (anonymous):

yeah sure... @_ashleeey

OpenStudy (anonymous):

A recent earthquake measured 6.8 on the Richter scale. How many times more intense was this earthquake than an earthquake that measured 4.7 on the Richter scale?

OpenStudy (anonymous):

hi. @_ashleeey

OpenStudy (anonymous):

do you have any idea of this problem??

OpenStudy (anonymous):

ummmmm, im guessing some multiplication?

OpenStudy (anonymous):

No i mean do you know how the Ritcher scale works, the formula relating its magnitude and intensity??

OpenStudy (anonymous):

Nopee

OpenStudy (anonymous):

read this then at first, or else the solution I will provide, will be of no use http://www.sosmath.com/algebra/logs/log5/log56/log56.html

OpenStudy (anonymous):

are you reading? .@_ashleeey

OpenStudy (anonymous):

i read it

OpenStudy (anonymous):

so so the formula is M= log (I/S) where M is the magnitue, I is the intensity and S is the intensity of a standard earth quake

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@_ashleeey I posted the solution in the attachment. read it and tell me if you have ay problem

OpenStudy (anonymous):

@Princer_Jones how come theres subtracting involved?

OpenStudy (anonymous):

we subtracted to eliminate S, which is unknown here